# BPINN-EM-Post: Stochastic Electromigration Damage Analysis in the Post-Void Phase based on Bayesian Physics-Informed Neural Network

### v1

@@ III-B Estimation of variations in EM stress
Here, $\bar{\sigma}$ represents the EM stress outcomes calculated using the parameter set $\mathbf{x}$ with the addition of noise $v$ from variability. $\mathcal{F}(.)$ could be any solver employed to solve combination of equations (1) and  (LABEL:eq:korhonen_equation_single_wire_postvoid), such as the FEM-based COMSOL method [39], or the FDM-based EMSpice method [18]. Variation noise is typically assumed to be independent and normally distributed with a mean of zero and a specified variance $Var_{\sigma}$. A common technique for estimating the mean and variance in EM stress calculations, as described by combinations of equations  (1) and  (LABEL:eq:korhonen_equation_single_wire_postvoid), involves direct Monte Carlo simulation. Nonetheless, Monte Carlo simulations necessitate repeated runs of combination of equations (1) and  (LABEL:eq:korhonen_equation_single_wire_postvoid) for multisegment interconnects, which can become exceedingly costly, particularly as the number of segments in the interconnect structure grows.
@@
Hence, in this work BPINN-EM-Post, we are combining fast and accurate analytical solutions with Bayesian PINN to efficiently estimate the variations in the EM stress on multisegment interconnect structure. This hierarchical combination of the analytical solution and Bayesian PINN provides accurate estimation of variations in EM stress significantly faster than the conventional Monte Carlo simulations.
IV Proposed BPINN-EM-Post
Figure 1: Framework of proposed BPINN-EM-Post variational EM simulator
IV-A Prior for BPINN
@@
As shown in Fig 1, a fully connected network serves as a surrogate model for our Bayesian physics-informed neural network (BPINN). Let each layer of the network be denoted as $l=1,2,3,\dots,L$ (where $L\geq 1$ is the number of layers). The output of the $l^{th}$ layer, $\mathbf{z}_{l}\in\mathbb{R}^{N_{l}}$, is expressed as:
$\mathbf{z}_{l}=\phi\left(\mathbf{w}_{l-1}\mathbf{z}_{l-1}+\mathbf{b}_{l-1}\right),$ (4)
@@
where $\phi$ is the non-linear activation function, $\mathbf{w}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ represents weight matrices, and $\mathbf{b}\in\mathbb{R}^{N_{l+1}}$ denotes bias vectors. The network’s output is given by:
$\tilde{\mathbf{F}}(\mathbf{x};\omega)=\mathbf{w}_{L}\mathbf{z}_{L}+\mathbf{b}_{L},$ (5)
@@
where $\mathbf{x}$ is the input vector. In this work, the Bayesian network processes information at all inter-segment junctions within a multisegment interconnect structure. The input is $\mathbf{x}=\{T,J_{L},J_{U},J_{R},J_{D}\}$, where $T$ is the temporal vector (time information) and $J_{L},J_{U},J_{R},J_{D}$ are current densities for segments connected to the left, upper, right, and downward positions. The output $\tilde{\mathbf{F}}$ is 3-dimensional, with components $F_{U}$, $F_{R}$, and $F_{D}$ representing flux through the upper, right, and downward segments.
@@
The unknown parameters $\omega$ comprise all weight matrices $\mathbf{w}$ and bias vectors $\mathbf{b}$. For the Bayesian neural network, $\omega$ is assigned a prior such that its components are independent Gaussian distributions with mean 00 and variances $\mathbf{Var}_{w,l}$ and $\mathbf{Var}_{b,l}$ for all layers [40]. As the hidden layer width approaches infinity, $\tilde{\mathbf{F}}(\mathbf{x})$ converges to a Gaussian process [40].
IV-B Generating observations using analytical solutions
@@
To generate stress solutions $\tilde{\sigma}(x,t;\omega)$ using $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, we apply analytical solutions proposed in [28]. Before using the output from the Bayesian network, $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, in the analytical solution, we perform linear transformations on it.
@@
At any inter-segment junction of a multisegment interconnect structure, up to four segments may connect. The three components of $\tilde{\mathbf{F}}$ represent flux information, which we denote as $F_{U}$, $F_{R}$, and $F_{D}$, corresponding to the upper, right, and downward segments as shown in Fig 1. A linear transformation is applied to compute the flux through the remaining segment. For example, flux through the left segment is calculated as $F_{L}=F_{U}+F_{R}+F_{D}$, enforcing atomic flux conservation, as described in Eq (1).
@@
As discussed in Section III, two types of segments exist in the post-voiding phase of a multisegment interconnect structure: one segment containing a void (Fig 2(b)) and voidless segments (Fig 2(a)). Given $F^{-}$ and $F^{+}$, stress $\tilde{\sigma}(x,t)$ at any point and time on a single wire segment can be accurately calculated.
(a) A single segment without void.
(b) A single segment with void at $x=L$.
Figure 2: Single interconnect segments that are part of a multisegment interconnect tree in the postvoiding phase.
For analytical solutions, we define the stress gradient variables at the left or bottom node as $\phi^{-}(t)$ and at the right or top node as $\phi^{+}(t)$. Consequently, Korhonen’s equation, modified with parametric boundary conditions for a wire segment without void, is given by
@@
$BC:\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{-}(t),x=0$
$BC:\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{+}(t),x=L$
@@
In Eq.(6), $\tilde{\sigma}(x,t)$ represents the revised stress response from the updated Korhonen’s equation, while $\phi^{-}(t)$ and $\phi^{+}(t)$ serve as tunable boundary conditions reflecting the stress gradients at the boundaries. It is assumed that $h(x)$ characterizes the initial stress state, corresponding to the stress at the nucleation moment $t_{nuc}$. The solution to (6) can be effectively derived using the Laplace transformation method, resulting in the following analytical solution for $\tilde{\sigma}(x,t)$ in the time domain:
$\hat{\sigma}(x,t)$ $=\sum_{p=0}^{P}\{-\frac{d\phi^{-}(t)}{dt}*\left(g(\zeta_{1}(p,x,L),t)+g(\zeta_{3}(p,x,L),t)\right)$ (7)
@@
$+\frac{2}{L}\sum_{m=1}^{M}(-1)^{m}\{e^{-M}\cos(m\pi x/L)$
×$\times\int_{0}^{L}h(u)\cos\left(m\pi(L-u)/L\right)du\}$

### v2

@@ II-B Estimation of Variations in EM Stress Evolution
here $\bar{\sigma}$ represents the EM stress outcomes calculated using the parameter set $\mathbf{x}$ with the addition of noise $v$ from variability. $\mathcal{F}(.)$ could be any solver employed to solve combination of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), such as the FEM-based COMSOL method [16], or the FDM-based EMSpice tool [13]. Variation noise is typically assumed to be independent and normally distributed with a mean of zero and a specified variance $Var_{\sigma}$ [28]. A common technique for estimating the mean and variance in EM stress calculations, as described by combinations of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), involves direct Monte Carlo simulation. Nonetheless, Monte Carlo simulations necessitate repeated runs of combination of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) for multi-segment interconnects, which can become exceedingly expensive, particularly as the number of segments in the interconnect structure grows.
The impact of variational or stochastic EM on the lifespan of wires has been explored in previous studies. [29] highlights that as technology advances, both the MTTF and the distribution of failure times deteriorate, posing ongoing challenges for advanced interconnects. Early work [1] introduced a method that incorporates variability into EM analysis. However, this approach examines a single wire segment individually and uses both analytical and numerical techniques to compute the time to failure. It considers two forms of variability: circuit-level variability, where the resistance of wires is treated as a random variable (RV), and wire geometry-level variability, where imperfections such as bumps and necking in wire geometries are RVs due to lithography-induced variations.  [30] investigates both global and local process variations in power grids, applying a Hermite polynomial chaos method for variational EM analysis and final lifetime estimation using Black’s EM model. In [31], a comprehensive analysis of the entire power grid is conducted, taking into account the inherent redundancy of grids by solving Korhonen’s PDE using finite difference techniques. Here, EM diffusivity is treated as a RV, and Monte Carlo approach is employed for the variational analysis, supplemented by ad-hoc acceleration techniques. [28] primarily focuses on variability in input currents, where stochastic current models are used to model currents in functional blocks, and variance or covariance calculations are done via matrix-solving for entire power grids, which remains computationally expensive. Work in [32] introduced a Bayesian PINN framework aimed at quantifying noise in observations and addressing the overfitting challenges typical of standard PINNs, and showed effectiveness on standard, relatively straightforward problems. However, this method applies Bayesian-based sampling to the entire PINN during training, which can be quite time-consuming, especially for large-scale domains such as multi-segment interconnect structures.
@@
Hence, in our proposed BPINN-EM-Post framework, we combine fast and accurate analytical solutions with Bayesian PINN to efficiently estimate the variations in the EM stress on multi-segment interconnect structures. This hierarchical combination of the analytical solution and Bayesian PINN provides accurate estimation of variations in EM stress significantly faster than the conventional Monte Carlo simulations. In this work we follow the problem setting in [28] that considers the stochastic EM stress caused by fluctuations of input current density. Other variations such as manufacture non-idealities can be modeled and analyzed through similar approach.
III Proposed BPINN-EM-Post Framework for Stochastic EM Analysis
III-A Estimation of Prior for BPINN
The overall framework of our proposed BPINN-EM-Post is shown in Fig. 1. A fully connected network serves as a surrogate model for our Bayesian physics-informed neural network (BPINN) framework, based on prior work [20, 24] and the demonstrated capability of Multi-layer Perceptron (MLP) architecture for this prediction task.
Figure 1: Framework of the proposed BPINN-EM-Post variational EM simulator
@@ III-A Estimation of Prior for BPINN
Let each layer of the network be denoted as $l=1,2,3,\dots,L$ (where $L\geq 1$ is the number of layers). The output of the $l^{th}$ layer, $\mathbf{z}_{l}\in\mathbb{R}^{N_{l}}$, is expressed as Eq. (4),
$\mathbf{z}_{l}=\phi\left(\mathbf{w}_{l-1}\mathbf{z}_{l-1}+\mathbf{b}_{l-1}\right)$ (4)
@@
where $\phi$ is the non-linear activation function, $\mathbf{w}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ represents weight matrices, and $\mathbf{b}\in\mathbb{R}^{N_{l+1}}$ denotes bias vectors. The network’s output is given by Eq. (5),
$\tilde{\mathbf{F}}(\mathbf{x};\omega)=\mathbf{w}_{L}\mathbf{z}_{L}+\mathbf{b}_{L}$ (5)
where $\mathbf{x}$ is the input vector.
@@
In this work, the Bayesian network processes information at all inter-segment junctions within a multi-segment interconnect structure. The input is $\mathbf{x}=\{T,J_{L},J_{U},J_{R},J_{D}\}$, where $T$ is the temporal vector (time information) and $J_{L},J_{U},J_{R},J_{D}$ are current densities for segments connected to the left, upper, right, and downward positions. The output $\tilde{\mathbf{F}}$ is 3-dimensional, with components $F_{U}$, $F_{R}$, and $F_{D}$ representing flux through the upper, right, and downward segments.
@@
The unknown parameters $\omega$ comprise all weight matrices $\mathbf{w}$ and bias vectors $\mathbf{b}$. For the BNN, $\omega$ is assigned a prior such that its components are independent Gaussian distributions with mean 0 and variances $\mathbf{Var}_{w,l}$ and $\mathbf{Var}_{b,l}$ for all layers [33]. As the hidden layer width approaches infinity, $\tilde{\mathbf{F}}(\mathbf{x})$ converges to a Gaussian process [33].
III-B Generating Observations with Analytical Solutions
@@
To generate stress solutions $\tilde{\sigma}(x,t;\omega)$ using $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, we adapt the analytical solutions proposed in [22]. Before using the output from the Bayesian network, $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, in the analytical solution, we perform linear transformations on it.
@@
At any inter-segment junction of a multi-segment interconnect structure, up to four segments may connect. The three components of $\tilde{\mathbf{F}}$ represent flux information, which we denote as $F_{U}$, $F_{R}$, and $F_{D}$, corresponding to the upper, right, and downward segments as shown in Fig. 1. A linear transformation is applied to compute the flux through the remaining segment. For example, flux through the left segment is calculated as $F_{L}=F_{U}+F_{R}+F_{D}$, enforcing atomic flux conservation, as described in Eq. (1).
@@
As discussed in Sec. II, two types of segments exist in the post-voiding phase of a multi-segment interconnect structure: a voidless segments as in Fig. 2(a), and a segment containing a void at terminal as shown in Fig. 2(b). Given $F^{-}$ and $F^{+}$, stress $\tilde{\sigma}(x,t)$ at any location and time on a single wire segment can be accurately calculated.
For analytical solutions, we define the stress gradient variables at the left or bottom node as $\phi^{-}(t)$ and at the right or top node as $\phi^{+}(t)$. Consequently, Korhonen’s equation, modified with parametric BCs for wire segment without void is given in Eq. (6),
@@
$BC$ $:\frac{\partial\tilde{\sigma}(x,t)}{\partial x}\bigg|_{x=0}=\phi^{-}(t),$
$BC$ $:\frac{\partial\tilde{\sigma}(x,t)}{\partial x}\bigg|_{x=L}=\phi^{+}(t),$
@@
((a)) Segment without void.
((b)) Segment with void at $x=L$.
Figure 2: Single interconnect segments that are part of a multi-segment interconnect tree in the post-voiding phase.
In Eq. (6), $\tilde{\sigma}(x,t)$ represents the revised stress response from the updated Korhonen’s equation, while $\phi^{-}(t)$ and $\phi^{+}(t)$ serve as tunable BCs reflecting the stress gradients at the boundaries. It is assumed that $h(x)$ characterizes the initial stress state, corresponding to the stress at the nucleation moment $t_{nuc}$. The solution to Eq. (6) can be effectively derived using the Laplace transformation method, resulting in the following analytical solution for $\tilde{\sigma}(x,t)$ in the time domain, as shown in Eq. (7),
$\tilde{\sigma}(x,t)$ $=\sum_{p=0}^{P}\{-\frac{d\phi^{-}(t)}{dt}*\left(g(\zeta_{1}(p,x,L),t)+g(\zeta_{3}(p,x,L),t)\right)$ (7)
@@
$+\frac{2}{L}\sum_{m=1}^{M}(-1)^{m}\{e^{-M}\cos(\frac{m\pi x}{L})$
$\times\int_{0}^{L}h(u)\cos\left(\frac{m\pi(L-u)}{L}\right)du\}$
In this context, $p$ and $m$ are positive integers, with $0\leq p\leq P$ and $1\leq m\leq M$. It is recommended to maintain the three predominant components by setting $P=2$ [21]. The value of $M$ is established through experimental methods and is elaborately discussed in the results section. In Eq. (7), the convolution operator is denoted by $\ast$, which is defined as $a(t)\ast b(t)=\int_{0}^{t}a(\tau)b(t-\tau)d\tau$. This convolution is efficiently executed using the Gauss-Legendre quadrature algorithm [21]. The functions $\zeta_{1}(p,x,L),\ \zeta_{2}(p,x,L),\ \zeta_{3}(p,x,L)$, $\zeta_{4}(p,x,L),$ and $g(x,t)$ are specified in Eq. (8) as follows,
$\zeta_{1}(p,x,L)=(2p+2)L-x,\$(8)
$\zeta_{2}(p,x,L)=(2p+1)L-x,$
$\zeta_{3}(p,x,L)=(2p)L+x,\$
$\zeta_{4}(p,x,L)=(2p+1)L+x,$
$g(x,t)=2\sqrt{\frac{kt}{\pi}}e^{-\frac{x^{2}}{4kt}}-x\cdot\texttt{erfc}\{\frac{x}{2\sqrt{kt}}\}$
For the segments with void, we assume that one end of the segment could either be an inter-segment junction or a terminal boundary, while the other end corresponds to a void node. Fig. 2(b) illustrates a single segment with a void at the $x=L$ end. At $x=0$, which may represent either an inter-segment junction or a terminal boundary, a modifiable BC from the voidless scenario is applied. We modify the BC at $x=L$ to $\tilde{\sigma}(L,t)=0$. When using a very small $\delta$ in Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), the BC at the site of the void mirrors the approach in [34]. The BCs for the segment with a void at the subsequent node of a wire segment, i.e., at $x=L$, are expressed as Eq. (9),
$BC:\frac{\partial\tilde{\sigma}(x,t)}{\partial x}\bigg|_{x=0}=\phi^{-}(t)$ (9)
$BC:\tilde{\sigma}(x,t)|_{x=L}=0$
Similar to the solution for the segment without a void, the solution for the single segment with a void location at $x=L$ is derived using the Laplace transform method as shown below in Eq. (10),
$\tilde{\sigma}(x,t)=$ $\sum_{p=0}^{P}(-1)^{p}\{\frac{d\phi^{-}(t)}{dt}*\left(g(\zeta_{1}(p,x,L),t)-g(\zeta_{3}(p,x,L),t)\right)$ (10)
$+\frac{2}{L}\sum_{m=1}^{M}(-1)^{m+1}\{e^{-M}\cos(\frac{(m-0.5)\pi x}{L})$
$\times\int_{0}^{L}h(u)\sin\left(\frac{(m-0.5)\pi(L-u)}{L}\right)du\}$
Similarly, when the void is positioned at the preceding (left/bottom) end, i.e., at $x=0$, the BC is defined as $\tilde{\sigma}(x,t)=0$ at $x=0$ and $\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{+}(t)$ at $x=L$. Using the same Laplace transform method, the analytical solution for this scenario can also be derived. In Eq. (7) and Eq. (10) the values $\phi^{-}(0)$ and $\phi^{+}(0)$ can be calculated using the current density information at each inter-segment junctions [22, 21]. For each single segment, we use the flux information from Bayesian network as $\frac{d\phi^{-}(t)}{dt}=F^{-}(\mathbf{x};\omega)$ and $\frac{d\phi^{+}(t)}{dt}=F^{+}(\mathbf{x};\omega)$.
The analytical solutions in Eq. (7) and Eq. (10) estimate the stress distribution for single segments within a multi-segment interconnect structure, given the flux information. For accurate stress distribution, all relevant physical conditions must be satisfied. While calculating the flux information, atomic flux conservation was enforced. Additionally, as per Eq. (1), stress continuity must be maintained at inter-segment junctions. This implies that the stress in all segments connected to a junction should be equal.
To ensure stress continuity, we optimize the BPINN using the stress continuity condition $\mathcal{L}$, which is defined in Eq. (11),
$\mathcal{L}=\frac{1}{N_{I}\times K_{i}}\sum_{i=1}^{N_{I}}\sum_{k=2}^{K_{i}}\left(\tilde{\sigma}_{k}(x_{\sigma};\omega)-\tilde{\sigma}_{k-1}(x_{\sigma};\omega)\right)^{2}$ (11)
where $\tilde{\sigma}_{k}(\omega)$ represents the EM stress at the boundary of the $k^{th}$ segment at inter-segment junction $i$, $K_{i}$ is the number of segments connected to junction $i$, and $N_{I}$ is the total number of inter-segment junctions in the multi-segment structure. Since $\tilde{\sigma}$ depends on $F^{-}(\mathbf{x};\omega)$ and $F^{+}(\mathbf{x};\omega)$, it is expressed as a function of $\omega$ and the parameters influencing stress, denoted as $x_{\sigma}$, for simplicity.
III-C Evaluation of Posterior for BPINN
In this work, the Bayesian network takes time and current information at each inter-segment junction of the multi-segment interconnect structure as input, i.e., $\mathbf{x}=\{T,J_{L},J_{U},J_{R},J_{D}\}$. Using the flux information from the network’s output, we enforce atomic flux conservation to calculate the flux values $F^{-}(\mathbf{x};\omega)$ and $F^{+}(\mathbf{x};\omega)$ at the endpoints of each segment.
Therefore, for this case we can define the dataset as Eq. (12),
here we assume that the observations are variational, and can be represented as in Eq. (13),
$\bar{\mathcal{L}}=\mathcal{L}(x_{\mathcal{L}}^{(i)})+\epsilon^{(i)},i=1,2,3,..N^{\mathcal{L}}$ (13)
where $\epsilon$ is independent Gaussian noise with zero mean and variance of $\mathbf{Var_{\mathcal{L}}}$. The likelihood of the observations can be calculated as Eq. (14),
$P(\mathcal{D}|\omega)=$ $\prod_{i=1}^{N_{\mathcal{L}}}\frac{1}{\sqrt{2\pi}Var_{\mathcal{L}}^{(i)}}\times exp\left[-\frac{\left(\mathcal{L}\left(x_{\mathcal{L}}^{(i)};w\right)-\bar{\mathcal{L}}^{(i)}\right)^{2}}{2Var_{\mathcal{L}}^{(i)}}\right]$ (14)
Utilizing this likelihood, the posterior can be obtained as,
here the symbol $\propto$ signifies ’equality up to a constant’. The probability of the dataset $P(D)$ is not readily solvable through analytical methods. Therefore, in practical scenarios, we only obtain an un-normalized expression $P(\omega|\mathcal{D})$.
This work BPINN-EM-Post employs the Hamiltonian Monte Carlo (HMC) method to navigate the parameter space of our BNN model. HMC is a gradient-based Markov Chain Monte Carlo (MCMC) technique that utilizes Hamiltonian dynamics for parameter exploration. Our HMC procedure begins by simulating Hamiltonian dynamics through a numerical integration method, followed by a correction using the Metropolis-Hastings acceptance step. Given a dataset $\mathcal{D}$, we postulate the target posterior distribution for $\omega$ as shown in Eq. (16),
$P(\omega|\mathcal{D})\simeq exp(-U(\omega))$ (16)
where $U(\omega)=-\log(P(\mathcal{D}|\omega))-\log(P(\omega))$ serves as the potential energy in the system. To sample from the posterior distribution, HMC introduces an auxiliary momentum variable $r$, which is used to formulate a Hamiltonian system as,
$H(\omega,r)=U(\omega)+\frac{1}{2}r^{T}\mathrm{M}^{-1}r$ (17)
Here, $M$ is a mass matrix usually set as identity matrix. HMC then generates samples from the joint distribution of $(\omega,r)$ as,
$\pi(\omega,r)\sim exp(-U(\omega)-\frac{1}{2}r^{T}\mathrm{M}^{-1}r)$ (18)
where the samples of $\omega$ have a marginal distribution as we discard the samples of $r$. These samples are derived from the following Hamiltonian dynamics as in Eq. (19),
$d\omega=\mathrm{M}^{-1}rdt,\ \ dr=-\nabla U(\omega)dt$ (19)
Eq. (19) is discretized through the leapfrog method, and the Metropolis-Hastings step is employed to minimize the discretization error. More implementation details on the HMC algorithm are explained in [33].
III-D Estimation of Variations in EM Stress
Using HMC, we sample the optimal Bayesian network parameters, $\omega$, to generate samples of flux information, $\small\left\{\tilde{\mathbf{F}}(\mathbf{x};\omega^{(i)})\right\}_{i=1}^{M}$. From these, we derive samples of $\{\mathbf{F}^{-}\}$ and $\{\mathbf{F}^{+}\}$. For each wire segment, these flux samples are used to compute $M$ stress solution samples, $\{\tilde{\sigma}(x,t)\}$.
Notably, when obtaining the optimal parameter samples via HMC, we utilize only stress observations at the boundaries of single wire segments within the multi-segment interconnect structure. This significantly reduces the variables in the loss function $\mathcal{L}$, enabling efficient training. Once the framework is fully trained and optimal parameters are obtained, we calculate the flux information at all wire segment boundaries. This flux information is termed optimal as it satisfies flux conservation and stress continuity at inter-segment junctions. Using the optimal flux information samples, we compute accurate stress distribution samples for individual segments through analytical solutions. Additionally, variations in stress solutions, such as mean and standard deviation, are estimated from the calculated stress samples.
IV Experimental Results
IV-A Experimental Setup
The proposed BPINN-EM-Post framework is fully developed using Python/PyTorch. The training and test procedures are conducted on a Linux server equipped with two Intel 22-core E5-2699 CPUs, 320 GB of memory, and an Nvidia TITAN RTX GPU. A three-layer MLP with layer structure [I5-FC50-FC50-O3] is used in BPINN flux predictor.
IV-B Data Preparation and Scaling
To determine the stress distribution in a multi-segment interconnect structure, we first gather the physical and material attributes of the interconnect segments. Each segment $i$ is characterized by its physical properties, including length $L_{i}$ and width $w_{i}$, while assuming uniform thermal conductivity $\kappa$ across all segments. Spatial points $x$ are sampled within $[0,L]$, and temporal points are chosen from $t=0$ to $1\times 10^{9}$ s at intervals of $1\times 10^{7}$ s. In this analysis, we incorporate the nucleation-phase solutions for these interconnects that are easily pre-established by existing methods [13, 16, 24] consistent to this work, providing initial stress distribution $h(x)$ and void locations.
((a)) Single segment without void.
((b)) Single segment with void at left node.
((c)) Single segment with void at right node.
Figure 3: Comparison of the stress distribution obtained from analytical solution with the stress distribution from FEM-based COMSOL for single interconnect segment for three different cases.
TABLE I: Performance and accuracy comparisons with existing methods [16][13]
# of wires COMSOL [16] EMSpice [13] The Proposed BPINN-EM-Post
Runtime Runtime Error v.s. COMSOL Error v.s. EMSpice Training time Inference time Speedup v.s. COMSOL Speedup v.s. EMSpice